A body having initial kinetic energy 2J collides with identical body at rest. The maximum loss of kinetic energy found in the collision is?

Grade:12th

2 Answers

Adarsh
6 years ago
Maximum loss of kinetic energy is possible when collision is perfectly inelastic when both body stick together and move with a velocity .. The loss of kinetic energy in inelastic collision is half the kinetic energy before collision .. 
So the loss in Kinetic energy will be 1j ...
Khimraj
6 years ago
let mass of body is m, initial velocity is u and final velocity is v.
As given ( ½)mu2 = 2J.
Maximum loss will be when collision is inelastic.
So final velocity of combined mass will be half of initial.
So final kinetic energy will be 2*( ½)*m*(u/2)2 =mu2/4 = 1J.
So loss in kinetic energy = 2J – 1J = 1J.
Hope it clears.